A current is forced through a cell consisting of two platinum electrodes in which each is immersed in a solution of CuSO4 and H2SO4. During the process copper is deposited at the cathode while evolution of oxygen occurs at the anode. The standard potential (Eo) for the reaction occurring at cathode and anode are 0.34V and 1.23V respectively.
(i) Write down the chemical equation for both the cathodic and anodic reactions.
(ii) What is the overall reaction and determine the theoretical decomposition potential?
(iii) Estimate the potential required for the operation of the cell if an additional 0.5V is applied to overcome the resistance of the cell to the current.
1
Expert's answer
2020-09-22T03:35:25-0400

( i ) Cathode = Cu2+ + 2e = Cu .

Anode = 2OH- = H2O + 1/ 2 O2 + 2e .


( ii ) Over all reaction

Cu2+ + 2OH- = Cu + H2O + 1/ 2 O2 .


( iii) Electrode Potential =( 0.34 + 1.23 ) V = 1.57 V .


 the potential required  = 1.57 V Answer .